# Learning Effective Representations for Retrieval Using Self-Distillation with Adaptive Relevance Margins

### v1

@@ Abstract.
Representation-based retrieval models, so-called biencoders, estimate the relevance of a document to a query by calculating the similarity of their respective embeddings. Current state-of-the-art biencoders are trained using an expensive training regime involving knowledge distillation from a teacher model and batch-sampling. Instead of relying on a teacher model, we contribute a novel parameter-free loss function for self-supervision that exploits the pre-trained language modeling capabilities of the encoder model as a training signal, eliminating the need for batch sampling by performing implicit hard negative mining. We investigate the capabilities of our proposed approach through extensive ablation studies, demonstrating that self-distillation can match the effectiveness of teacher distillation using only 13.5% of the data, while offering a speedup in training time between 3x and 15x compared to parametrized losses. Code and data is made openly available.111github.com/webis-de/adaptive-relevance-margin-loss
@@ 1. Introduction
In information retrieval, transformer-based bi-encoders are used as effective retrieval models. They estimate a document’s relevance to a query by computing the similarity of their embeddings. Queries and documents are represented independently, which allows for pre-computing and indexing all document representations offline and computing only the query representation at retrieval time. bi-encoders are thus appealing in practice since they can be easily scaled. In contrast, transformer-based cross-encoders are a type of model that estimates relevance jointly by computing a score for each document-query pair at retrieval time, achieving higher effectiveness than bi-encoders at the expense of a higher computational cost and inference latency. They are thus usally reserved for multi-stage reranking, such that only the top-k documents of an initial ranking produced by a more efficient bi-encoder are re-ranked.However, highly effective bi-encoders come at the cost of a highly expensive training regime based on knowledge distillation. Figure 1 (left) illustrates this process, where the relevance of each training sample is first estimated by a cross-encoder teacher model. The estimated relevance scores are first used for a computationally intensive batch sampling process, where an optimal distribution of relevance scores within each batch is ensured according to the teacher model, which then provides supervision for training a bi-encoder student model with a margin-based loss function. While this knowledge distillation setup yields highly effective bi-encoder ranking models, we highlight three drawbacks: (1) A teacher model must be available for the desired dataset and domain. cross-encoder models are usually chosen as teacher models to maximize effectiveness (Hofstätter et al., 2020). Training a cross-encoder model for knowledge distillation is computationally expensive and may not be possible for domains with little ground-truth ranking data. (2) Teacher scores must be inferred for every training sample, requiring a forward pass. This adds inference cost onto a process that is already prohibitively expensive. Improving training efficiency is crucial for information retrieval, as the majority of energy spent over the lifetime of a model, especially models that are efficient at retrieval time, goes into training (Scells et al., 2022). (3) Batch sampling prohibits continuous learning on new data. The current best batch sampling procedures rely on clustering the training data set (Hofstätter et al., 2020). A static training dataset is thus required, rendering it challenging to adapt such approaches in settings where training data is collected continuously, such as online learning-to-rank.
In this paper, we propose a novel training regime and loss function for bi-encoders (Figure 1, right) that eliminates the complexities of previous approaches while maintaining competitive effectiveness. Instead of using a teacher model, we leverage the pre-trained text similarity capabilities of the encoder model to provide a supervision signal in a self-distillation setup. This eliminates the need for expensive exhaustive inference over the training dataset, does not rely on batch sampling techniques, and is highly data efficient. We then conduct systematic ablation studies which demonstrate that self-distillation can match the effectiveness of teacher-distillation based training regimes, with nDCG@10 and Recall@1000 scores of no significant difference, while using only 13.5% of the data. Finally, our proposed loss function is hyperparameter free, eliminating the need for expensive hyperparameter tuning.
@@ 2. Preliminaries & Related Work
We define as a set of queries representing user information needs and as a set of documents potentially containing relevant information. The objective of a ranking function is to order the documents based on their relevance to each query , accomplished with a relevance model , which assigns relevance scores that establish an order. We summarize related work on operationalizing with representation learning, covering encoder models (Section 2.1), knowledge distillation (Section 2.2), loss functions (Section 2.3), and negative mining (Section 2.4).
@@ 2.1. Relevance Estimation with Encoder Models
Text encoder models like BERT (Devlin et al., 2019) operationalize in two main ways: bi-encoders and cross-encoders (Zhao et al., 2022). Bi-encoders independently embed the query and each document using an encoder model , typically by using the representation of the special [CLS] token (optionally processed by a linear layer) (Chang et al., 2020; Gao and Callan, 2021), and express relevance in terms of a similarity function as (Karpukhin et al., 2020; Chang et al., 2020; Gao et al., 2021). In contrast, cross-encoders process both query and document jointly, operationalizing as a relevance classification of the joint representation (Nogueira et al., 2019): , where is string concatenation and is a binary classifier assigning a probability. While bi-encoders are more efficient (Zhao et al., 2022), cross-encoders capture complex interactions between query and document (Nogueira et al., 2019), yielding higher effectiveness albeit with higher latency at query time, as document representations cannot be precomputed. We focus our investigation on training bi-encoders with comparable effectiveness to cross-encoders.
@@ 2.2. Knowledge Distillation
Knowledge distillation (Hinton et al., 2015) aims to train a student model to reproduce the output of a teacher model (Hu et al., 2023). It is commonly employed for model compression, i.e., to develop smaller and faster models capable of maintaining effectiveness at reduced computational effort (Gou et al., 2021). In retrieval tasks, distillation, as seen in works by Hofstätter et al. (Hofstätter et al., 2020, 2021), trains bi-encoder models using precomputed relevance scores from a cross-encoder teacher model. Although distillation can successfully impart ranking effectiveness to bi-encoders, the use of cross-encoder teachers imposes constraints on the training process due to the computational cost of pre-computing relevance scores for a large set of query-document pairs. Thus, while distillation produces highly efficient student models, it incurs high computational costs for training.Self-distillation is a subfield of knowledge distillation, where the student and teacher models are identical (Hahn and Choi, 2019), i.e., leveraging a model’s own output to provide a training signal for a refined version of itself. The success of this approach is attributed to the rich output distribution, capturing additional information about training examples (Liu et al., 2021). While self-distillation is predominantly applied when teacher and student are trained towards the same task, it can also be used in a transfer setting (Zhou et al., 2022). Here, the pretrained capabilities of the model on a general task are exploited to train a more refined version of itself on a specialized target task. Self-distillation reduces the required computational effort, as no additional teacher inference is needed. We take inspiration from prior work in the literature about self-distillation to apply to the task of ranking.
@@ 2.3. Bi-Encoder Loss Functions
In the typical training setup for bi-encoder models, data comes in the form of triplets , comprising a query, a positive document, and a negative document. Various loss functions have been proposed within this setup: for instance, triplet loss aims to adjust the representation space so that the positive document is closer to the query than the negative document by a certain margin (Schroff et al., 2015). Karpukhin et al. (2020) minimize the negative log-likelihood between a single positive example and multiple negative examples. Xiong et al. (2021) minimize a classification loss function like binary cross-entropy, hinge loss, or negative log-likelihood between scores predicted for pairs of positive and negative examples. All aim to rearrange representations to prioritize relevance, i.e., learn a representation space in which for a query, a relevant passage is closer than a non-relevant passage by a margin .In knowledge distillation contexts, teacher scores provide a source of supervision. The MarginMSE loss (Hofstätter et al., 2020) adopts the concept of optimizing the mean squared error between the relevance difference, i.e., , and a desired target . The value of is dictated individually for each triplet by its predicted teacher model output. This notion of introducing an ‘adaptive’ target value for each instance also has broader applications beyond distillation. For instance, Li et al. (2020) use class distances in an image classification problem as adaptive target, Ha and Blanz (2021) use differences in ground-truth labels as adaptive targets for learning to rank, and Sharma et al. (2024) use the similarity between augmented views of the same sample. We apply self-distillation to predict an adaptive target for each training triplet using the bi-encoder model itself.
@@ 2.4. In-Batch Negatives & Hard Negative Mining
To increase the efficiency of triplet-based losses, additional training instances can be formed by considering documents from other triplets in the same batch as supplementary negative examples for a given query-positive pair (Zhan et al., 2021). Yet, a major caveat is that rapidly learns to correctly project trivial triplets, i.e., triplets where and are easy to distinguish w.r.t. their relevancy. Consequently, identifying ‘hard’ triplets, where distinguishing between positive and negative documents is challenging, becomes crucial for learning a robust latent space (Zhan et al., 2021). Yet, in-batch negatives are likely topically unrelated to a given query, and thus have limited learning potential.To alleviate this issue, batch sampling techniques have been devised that select batches of triplets such that the positive and negative documents are hard to distinguish in relevance, both within each pair and across pairs of in-batch negatives (Cai et al., 2022). For retrieval, Hofstätter et al. (2021) introduce balanced topic-aware sampling (TAS-B), where triplets are chosen to ensure that the relevance differences of query-positive and query-negative pairs across all training pairs in a batch adhere to a certain distribution. All triplets in the training dataset are put into bins based on their margin, and then batches are constructed by uniformly selecting triplets from all bins. A reproduction study by Wang and Zuccon (2023) found that further effectiveness is possible by discarding bins with high relevance differences, removing ‘easy’ negatives from consideration. By predicting adaptive targets, we propose that self-distillation can be used to perform ‘implicit’ negative sampling based on the magnitude of the target without explicitly sampling batches.
@@ 3. Self-Distillation With Adaptive Relevance Margins
We formulate three criteria to simplify the training process and increase the training efficiency of bi-encoders for retrieval and address the drawbacks of teacher-distillation approaches outlined before: (1) no teacher model for target estimation: the loss function should only rely on the information contained in provided training triplets, (2) no batch sampling procedure: the loss function should be effective with randomly ordered data, (3) no hyperparameters: the loss function should be applicable without computationally intensive hyperparameter tuning. In this section, we introduce adaptive relevance margins, a novel self-distillation loss function for retrieval fine-tuning. We motivate it starting from how a margin loss is traditionally calculated, and deriving several variants for the inner term of the loss incrementally.
@@
The traditional margin loss for a triplet of query, positive (relevant) document, and negative (non-relevant) document contrasts the query relevance margin, i.e., the difference between the cosine similarity of a query and a positive document and the cosine similarity between query and a negative document, with a target value (Hofstätter et al., 2021):222We explicitly omit writing the representation function in our notation for brevity, therefore already signify query/document representations.
This variant optimizes for a latent space where positive and negative documents are separated by a value of at least w.r.t. their query similarity. However, must be tuned for the best effectiveness.
@@
Rather than using the same value as a target across all training instances, knowledge distillation allows one to derive per-document target scores using a teacher model (Hofstätter et al., 2021). We propose using the bi-encoder model itself as a source for this teacher signal in a self-distillation setup. As BERT-style encoders are pretrained with a masked token or sentence prediction objective, they exhibit strong capabilities for measuring text similarity (Devlin et al., 2019; Liu et al., 2019). We can, therefore, replace the with the similarity between positive and negative documents for a query, scaled to the 0-1 range:
Rescaling the target is warranted since we desire a latent space where positive and negative documents are orthogonal. This adaptive margin may initially seem counterintuitive as the model learns to predict a larger difference in relevance for similar texts. However, this can be framed from the perspective of hard negatives: if the positive and negative document are similar, i.e., , they are a ‘hard’ training instance. This larger margin leads to an increased contribution to the overall loss and, thus, higher impact on the model parameters compared to a dissimilar positive and negative document pair, i.e., . Adaptive targets thus implicitly perform hard negative mining without hyperparameters.
@@
So far, the loss calculation has used only a single triple. While naïve in-batch negative variants of the static and adaptive targets losses can be derived by replacing with multiple , this yields sub-par results if not combined with batch sampling techniques. Instead, we extend the adaptive targets approach to approximate this sampling by using multiple different target values for each individual training triple.
The intuition behind this approach is that the target now forms a distribution over the negative documents in the batch. Instead of comparing each triplet in the batch with a single target, we repeatedly compare the margin to the original negative with targets from all of the negatives in a batch. In other words, these repeated comparisons estimate the average scaled similarity of a positive document to (a sample of) the population of negative documents. If this similarity is high (), the sample contains documents that are hard to distinguish w.r.t. relevancy and thus contains rich information to establish a positive-negative class separation. If this similarity is low (), the sample contains documents that are easy to distinguish w.r.t relevancy. A high in-batch negative document similarity translates to a large margin to maximize the impact on the loss, and vice versa.While adaptive targets opt to use in-batch information to increase the count of total available training triples, distributed targets instead exploit the in-batch distribution to make a more accurate estimation of the target for each original triple.One can also consider the distributed targets approach as optimizing for the implicit mean document distance. The reason optimizing it implicitly works and optimizing it explicitly, i.e., by taking the mean scaled document similarity within the loss as target, does not work is because optimizing it explicitly repositions only in latent space rather than all as the implicit mean does.
@@
To form a proper loss function, the previously derived inner terms are optimized using mean squared error (MSE), which calculates the average squared difference of the relevance margin and target over all samples in a batch. A batch consists of a set of queries , a set of associated positive (relevant) documents , and a set of associated negative (non-relevant) documents , with batch size . For all variants not relying on in-batch information, the error is calculated as
@@
Summary.
Figure 2 illustrates the behaviour of all three loss variants. The position of a query, a positive document, and negative documents are visualized in cosine similarity space. Green elements indicate the target values and red arrows indicate the repositioning in latent space induced by the respective loss. We only illustrate repositioning of negative documents for simplicity; in practice, all three, query, positive, and negative could be simultaneously affected. For (a) static targets, a relevance margin of is forced (green lines); for (b) adaptive targets, individual relevance margins for each negative are estimated based on document similarity (green dots); and for (c) distributed targets, for a single negative, the distribution of the document similarities of all negatives (green area) is used as margin (green dot).

### v2

@@ Abstract.
Representation-based retrieval models, so-called bi-encoders, estimate the relevance of a document to a query by calculating the similarity of their respective embeddings. Current state-of-the-art bi-encoders are trained using an expensive training regime involving knowledge distillation from a teacher model and batch-sampling. Instead of relying on a teacher model, we contribute a novel parameter-free loss function for self-supervision that exploits the pre-trained language modeling capabilities of the encoder model as a training signal, eliminating the need for batch sampling by performing implicit hard negative mining. We investigate the capabilities of our proposed approach through extensive experiments, demonstrating that self-distillation can match the effectiveness of teacher distillation using only 13.5% of the data, while offering a speedup in training time between 3x and 15x compared to parametrized losses.All code and data is made openly available.111Code: github.com/webis-de/adaptive-relevance-margin-loss
Run files: zenodo.org/records/11197962
by
@@ 1. Introduction
In information retrieval, transformer-based bi-encoders are used as effective retrieval models. They estimate a document’s relevance to a query by computing the similarity of their embeddings. Queries and documents are represented independently, which allows for pre-computing and indexing all document representations offline and computing only the query representation at retrieval time. Bi-encoders are thus appealing in practice since they can be easily scaled. In contrast, transformer-based cross-encoders are a type of model that estimates relevance jointly by computing a score for each document–query pair at retrieval time, achieving higher effectiveness than bi-encoders at the expense of a higher computational cost and inference latency. They are thus usually reserved for multi-stage re-ranking, such that only the top- documents of an initial ranking produced by a more efficient bi-encoder are re-ranked.
However, highly effective bi-encoders come at the cost of a highly expensive training regime based on knowledge distillation. Figure 1 (left) illustrates this process, where the relevance of each training sample is first estimated by a cross-encoder teacher model. The estimated relevance scores are first used for a computationally intensive batch sampling process, where an optimal distribution of relevance scores within each batch is ensured according to the teacher model, which then provides supervision for training a bi-encoder student model with a margin-based loss function. While this knowledge distillation setup yields highly effective bi-encoder ranking models, we highlight three drawbacks: (1) A teacher model must be available for the desired dataset and domain. Cross-encoder models are usually chosen as teacher models to maximize effectiveness (Hofstätter et al., 2020). Training a cross-encoder model for knowledge distillation is computationally expensive and may not be possible for domains with little ground-truth ranking data. (2) Teacher scores must be inferred for every training sample, requiring a forward pass. This adds inference cost onto a process that is already prohibitively expensive. Improving training efficiency is crucial for information retrieval, as the majority of energy spent over the lifetime of a model, especially models that are efficient at retrieval time, goes into training (Scells et al., 2022). (3) Batch sampling prohibits continuous learning on new data. The current best batch sampling procedures rely on clustering the training data set (Hofstätter et al., 2020). A static training dataset is thus required, rendering it challenging to adapt such approaches in settings where training data is collected continuously, such as online learning-to-rank. In this paper, after a careful review of related work (Section 2), we propose a novel training regime and loss function for bi-encoders (Figure 1, right) that simplifies previous approaches while maintaining competitive effectiveness (Section 3). Instead of using a teacher model, we leverage the pre-trained text similarity capabilities of the encoder model to provide a supervision signal in a self-distillation setup. This eliminates the need for expensive exhaustive inference over the training dataset, does not rely on batch sampling techniques, and is highly data efficient. Furthermore, our proposed loss function is free of hyperparameters. We then conduct systematic experiments (Section 4) which demonstrate that self-distillation can match and even surpass the effectiveness of teacher distillation-based training regimes in in-domain settings, and still sufficiently generalize to zero-shot ranking tasks, while using only 13.5% of the data (Section 5).
@@ 2. Preliminaries & Related Work
We define as a set of queries representing user information needs and as a set of documents potentially containing relevant information. A retrieval model maps pairs of a document and a query to a relevance score, where a higher score indicates higher relevance of to . We summarize related work on operationalizing with representation learning, covering encoder models (Section 2.1), knowledge distillation (Section 2.2), loss functions (Section 2.3), and negative mining (Section 2.4).
@@ 2.1. Relevance Estimation with Encoder Models
Text encoder models like BERT (Devlin et al., 2019) operationalize in two ways: bi-encoders and cross-encoders (Zhao et al., 2022). Bi-encoders independently embed the query and each document using an encoder model , typically by using the representation of the special [CLS] token (optionally processed by a linear layer) (Chang et al., 2020; Gao and Callan, 2021). They measure relevance in terms of a similarity function as (Karpukhin et al., 2020; Chang et al., 2020; Gao et al., 2021). In contrast, cross-encoders jointly process query and document, operationalizing as a relevance classification of the joint representation (Nogueira et al., 2019): , where is string concatenation and is a classifier assigning a probability of relevance. While bi-encoders are more efficient (Zhao et al., 2022), cross-encoders capture complex interactions between query and document (Nogueira et al., 2019), yielding higher effectiveness albeit with higher latency at query time, as document representations cannot be pre-computed. We focus our investigation on training bi-encoders with comparable effectiveness to cross-encoders.
@@ 2.2. Knowledge Distillation
Knowledge distillation (Hinton et al., 2015) aims to train a student model to reproduce the output of a teacher model (Hu et al., 2023). It is commonly employed for model compression, i.e., to develop smaller and faster models capable of maintaining effectiveness at reduced computational effort (Gou et al., 2021). In retrieval tasks, distillation, as proposed by Hofstätter et al. (2020, 2021), trains bi-encoder models using pre-computed relevance scores from a cross-encoder teacher model. Although distillation can successfully impart ranking effectiveness to bi-encoders, the use of cross-encoder teachers imposes constraints on the training process due to the computational cost of pre-computing relevance scores for a large set of query–document pairs. Thus, while distillation produces highly efficient student models, it incurs high computational costs for training.Self-distillation is a subfield of knowledge distillation, where the student and teacher models are identical (Hahn and Choi, 2019), i.e., leveraging a model’s own output to provide a training signal for a refined version of itself. The success of this approach is attributed to the rich output distribution, capturing additional information about training examples (Liu et al., 2021). While self-distillation is predominantly applied when the teacher and student models are trained towards the same task, it can also be used in a transfer setting (Zhou et al., 2022). Here, the pre-trained general capabilities of the model are exploited to train a more refined version of itself on a specialized target task.This reduces the required computational effort, as no additional teacher inference is needed. We take inspiration from prior work in the literature about self-distillation to apply to the task of ranking.
@@ 2.3. Bi-Encoder Loss Functions
In the typical training setup for bi-encoder models, data comes in the form of triplets , comprising a query, a positive document , and a negative document . Various loss functions have been proposed within this setup: for instance, triplet loss aims to adjust the representation space so that the positive document is closer (i.e., more relevant) to the query than the negative document by a certain margin (Schroff et al., 2015). Karpukhin et al. (2020) minimize the negative log-likelihood between a single positive document and multiple negative documents. Xiong et al. (2021) minimize a classification loss function like binary cross-entropy, hinge loss, or negative log-likelihood between scores predicted for documents pairs. All aim to rearrange representations to prioritize relevance, i.e., to learn a representation space in which for a query, a relevant document is closer than a non-relevant document by a margin .In knowledge distillation contexts, teacher scores provide a source of supervision. The MarginMSE loss (Hofstätter et al., 2020) adopts the concept of optimizing the mean squared error between the relevance difference, i.e., , and a desired target . The value of is dictated individually for each triplet by its predicted teacher model output. This notion of introducing an ‘adaptive’ target value for each instance also has broader applications beyond distillation. For instance, Li et al. (2020) use class distances in an image classification problem as adaptive target, Ha and Blanz (2021) use differences in ground-truth labels as adaptive targets for learning-to-rank, and Sharma et al. (2024) use the similarity between augmented views of the same sample. We apply self-distillation to predict an adaptive target for each training triplet using the bi-encoder model itself.
@@ 2.4. In-Batch Negatives & Hard Negative Mining
To increase the efficiency of triplet-based losses, additional training instances can use other triplets in the same batch as supplementary negative documents for a given pair of query and positive document (Zhan et al., 2021). A major caveat is that rapidly learns to correctly project trivial triplets, i.e., triplets where and are easy to distinguish with respect to their relevance. Consequently, identifying ‘hard’ triplets, where distinguishing between positive and negative documents is challenging, becomes crucial for learning a robust latent space (Zhan et al., 2021). Yet, in-batch negatives are likely topically unrelated to a given query, and thus have limited learning potential.To alleviate this issue, batch sampling techniques have been devised that select batches of triplets such that the positive and negative documents are hard to distinguish in terms of their relevance to , both within each pair and across pairs of in-batch negatives (Cai et al., 2022). For retrieval, Hofstätter et al. (2021) introduce balanced topic-aware sampling (TAS-B), where triplets are chosen to ensure that the relevance differences of query–positive and query–negative pairs across all training pairs in a batch adhere to a certain distribution. All triplets in the training data are put into bins based on their margin, and then batches are constructed by uniformly selecting triplets from all bins. A reproduction study by Wang and Zuccon (2023) found that effectiveness can be further improved by discarding bins with high relevance differences, removing ‘easy’ negatives from consideration. By predicting adaptive targets, we show that self-distillation allows ‘implicit’ negative sampling based on the magnitude of the target without explicitly sampling batches.
@@ 3. Self-Distillation With Adaptive Relevance Margins
We formulate three criteria to simplify the training process, increase the training efficiency of bi-encoders for retrieval, and address the drawbacks of teacher distillation approaches outlined before: (1) no teacher model for target estimation: the loss function should only rely on the information contained in provided training triplets; (2) no batch sampling procedure: the loss function should be effective with randomly ordered data; (3) no hyperparameters: the loss function should be applicable without hyperparameter tuning. In this section, we introduce adaptive relevance margins, a novel self-distillation loss function for retrieval fine-tuning. We motivate it starting from how a margin loss is traditionally calculated, and derive several variants for the inner term of the loss.
@@
The traditional margin loss for the -th training triplet consisting of a query , a positive (i.e., relevant) document , and a negative (i.e., less or non-relevant) document contrasts the query–document relevance margin (Hofstätter et al., 2021). It is calculated as the difference between the cosine similarity of and and the cosine similarity of and , with a target value :222For brevity, we omit the encoder from the loss functions.
This loss optimizes for a latent space where positive and negative documents are separated by a value of at least in terms of their similarity toa query, where is a hyperparameter to be tuned.
@@
Rather than using the same value as a target across all training instances, knowledge distillation allows one to derive per-document target scores using a teacher model (Hofstätter et al., 2021). We propose using the bi-encoder model itself as a source for this teacher signal through self-distillation. As BERT-style encoders are pre-trained with a masked token or sentence prediction objective, they exhibit strong capabilities for measuring text similarity (Devlin et al., 2019; Liu et al., 2019). We can, thus, replace with the normalized similarity between positive and negative documents for a query, scaled to the range:
Rescaling the target is warranted, since we desire a latent space where positive and negative documents are orthogonal. In particular, we designed this adaptive margin to learn to predict a larger difference in relevance for similar documents rather than dissimilar ones. If the positive and negative documents are similar (i.e., ), they are a ‘hard’ training instance, which results in a higher contribution to the overall loss and therefore a higher impact on the model parameters compared to dissimilar documents (i.e., ). Adaptive targets thus implicitly perform hard negative mining without hyperparameters.
@@
So far, the loss calculation has used only a single triplet . While naïve in-batch negative variants of the static and adaptive targets losses can be derived by replacing with multiple , this yields sub-par results if not combined with batch sampling techniques. Instead, we extend the adaptive targets approach to approximate this sampling by using multiple different target values for each individual training triplet:
The intuition behind this approach is that the target now forms a distribution over the negative documents in the batch. Instead of comparing each triplet in the batch with a single target, we repeatedly compare the margin to the original negative with targets from all of the negatives in a batch. In other words, these repeated comparisons estimate the average scaled similarity of a positive document to (a sample of) the population of negative documents . If the similarity is high (), the sample contains documents that are hard to distinguish in terms of relevance and thus comprises rich information to establish a positive–negative class separation. If the similarity is low (), the sample contains documents that are easy to distinguish in terms of relevance. A high in-batch negative document similarity translates to a large margin to maximize the impact on the loss, and vice versa.While adaptive targets opt to use in-batch information to increase the count of total available training triplets, distributed targets instead exploit the in-batch distribution to make a more accurate estimation of the target for each original triplet. One can also consider the distributed targets approach as optimizing for the implicit mean document distance. The reason optimizing it implicitly works and optimizing it explicitly (i.e., by taking the mean scaled document similarity within the loss as target) does not work is because optimizing it explicitly repositions only in latent space rather than all as the implicit mean does.
@@
To form a proper loss function, the previously derived pointwise terms are optimized using mean squared error (MSE), which calculates the average squared difference of the relevance margin and target over all samples in a batch. A batch consists of a set of queries , a set of associated positive (relevant) documents , and a set of associated negative (less or non-relevant) documents , with batch size . For all variants not relying on in-batch information, the error is calculated as
@@
Discussion.
Figure 2 illustrates the behavior of all three loss variants. The position of a query , a positive document , and a set of negative documents are visualized in cosine similarity space. Green elements indicate the target values and red arrows indicate the repositioning in latent space induced by the respective loss. We only illustrate repositioning of negative documents, for simplicity; in practice, all three, query, positive, and negative could be simultaneously affected. For (a) static targets, a relevance margin of is forced (green lines); for (b) adaptive targets, individual relevance margins for each negative are estimated based on document similarity (green dots); and for (c) distributed targets, for a single negative , the distribution of the document similarities of all negatives (green area) is used as margin (green dot).